OpenStudy (anonymous):

how would i simplify square root of 75x^2 multiplied by square root of 3^3?

OpenStudy (anonymous):

\[\sqrt{75x^2}\times \sqrt{3^3}\]?

OpenStudy (anonymous):

Yes that is the expression I need to simplify

OpenStudy (anonymous):

\[\sqrt{75x^2\times 3^3}\] \[\sqrt{25\times 3\times x^2\times 3^3}\]\[\sqrt{25\times 81\times x^2}\] \[5\times 9\times x\] \[45x\]

OpenStudy (anonymous):

assuming x is positive. otherwise you have to write \[45|x|\]

OpenStudy (anonymous):

so that is the simplified answer to that expression?

OpenStudy (anonymous):

yes there is nothing under the square root sign

OpenStudy (anonymous):

ok thank you very much!